you have $50 to spend on a party. Pizzas cost $10 each, ships cost $2 per bag , and soda costs $1.25 per bottle. write and equation in three variables for the number of pizzas, bags of chips, and bottles of soda you buy. find the intercept of the equation and graph the equation
Answer by Edwin McCravy(8909)

you have $50 to spend on a party. Pizzas cost $10 each, ships cost $2 per bag , and soda costs $1.25 per bottle. write and equation in three variables for the number of pizzas, bags of chips, and bottles of soda you buy. find the intercept of the equation and graph the equation
10x + 2y + 1.25z = 10

To find the x-intercept, substitute 0 for y and z and
solve for x


10x + 2(0) + 1.25(0) = 10
                 10x = 10
                   x = 1

So the x-intercept is the point (1,0,0)

To find the y-intercept, substitute 0 for x and z and
solve for y


10(0) + 2y + 1.25(0) = 10
                  2y = 10
                   y = 5

So the y-intercept is the point (0,5,0)

To find the z-intercept, substitute 0 for x and y and
solve for z


10(0) + 2(0) + 1.25z = 10
               1.25z = 10
                   z = 8

So the x-intercept is the point (0,0,8)

Sorry, we can't draw 3D graphs on here. But it is a plane
which passes through those three points.

Edwin